# Versatile Demonstration Interface: Toward More Flexible Robot Demonstration Collection

### v1

@@ IV-B Evaluation with Manufacturing Experts
After completing the three conditions of the rolling task, the participants completed the press-fitting task. Participants were instructed about the goals of the task and shown how to complete the task without the robot. Participants were prompted that they could switch demonstration modes at any point during the task, but that they needed to use each mode to complete at least one press fitting. Following the two tasks, participants engaged in a semi-structured interview where they were asked about their preferences, design feedback, and potential use cases for the prototype interface.
Metrics – After each condition of the rolling task, participants filled out the NASA Task Load Index (TLX) [49], the Systems Usability Scale (SUS) Questionnaire [50], and the Usefulness and Ease of Use 7-point subscales of the USE Questionnaire [51]. Given our small sample size, we report only descriptive statistics for the questionnaire responses. Responses from the semi-structured interview were synthesized into the key themes presented in the next section.
@@ V Results and General Discussion
Here we discuss the results of our user study, limitations that we plan to address in future work, and general conclusions. All participants were generally able to complete the rolling task with each demonstration modality. In the press-fitting task, all participants were able to complete the task using the VDI tool. The questionnaire response results are shown in Figure 4 and participants ranked preferences are shown in Table II. Two participants responses are omitted from the questionnaire results due to missed survey items.
@@ V-A User Study Interactions and Implications
The clear winner: natural demonstrations – The majority of participants ranked the natural mode as their top preference. Participants commented that it was the fastest way to accomplish the tasks [P7], was valuable for delicate motions [P4], and that they felt they could perform the task better compared to the other demonstration modes [P5, P8]. This was also reflected in the survey responses, where participants tended to rank the natural mode as lower task load and higher ease of use compared to the other two methods. Several participants attributed the benefits of this mode to its innate familiarity, including that it felt similar to using a pencil [P3] or a manual tool [P1]. For example, P2 commented “I am so used to working with my hands anyway, that it was a logical way to do things.” The most common complaint among participants, including those who didn’t prefer the natural mode, was the tracking accuracy. In the prototype, participants found that the range was very limited [P6] and that it could at times be difficult to guide the robot to certain views of the task. Given such limitations in range, P3 mentioned the natural mode may be most effective for 2D (e.g., tabletop) tasks where a similar camera angle can be used to track the hand tool longitudinally. Many participants commented that improvements to the tracking would increase their preference for the natural mode.
@@
A time and place for each demonstration mode – While the natural mode was generally preferred, participants acknowledged the value of accessing the three demonstration modes through a single interface, both to cater to personal preferences and to task characteristics. As P7 commented, “With the engineers that I work with… we have a broad range of right-out-of-school to thirty years’ experience. People would pick different methods for sure.” Generally, the natural method was preferred for tasks involving complex and dexterous motions. For teleoperation, participants saw value for tasks where the demonstrator couldn’t be safely near the robot, such as environments with bio-hazards [P6] or heavy payloads [P7]. Some participants also commented on the value of using teleoperation to scale down human input for small, precise motions (e.g., soldering) [P4, P9] or to avoid the fatigue associated with more manual demonstration methods of repetitive tasks [P4]. For kinesthetic teaching, participants thought it would be valuable for tasks where more direct control was desired, but where the robot could assist with pushing/pulling [P5], or maintaining orientation [P1, P2]. P8 commented kinesthetic teaching was a good compromise between precision and ease of use “where I didn’t have to do too much, but I had some control.”
Participants also expressed some value in switching modes within tasks. Several participants mentioned using one mode for more coarse motions (e.g., kinesthetic teaching) and one mode for more precise motions (e.g., natural or teleoperation) [P2, P3, P9]. However, P1 commented that it may not be valuable to switch from teleoperation often if it is primarily used for tasks where the operator is removed from the robot.
Outlook and Improvements – Participants were generally optimistic about a range of repetitive tasks that could be trained using a versatile demonstration interface; including assembly (e.g., pick and place, pressing fittings, wire assembly) [P1, P4, P5, P7, P9], stoppering [P1], grinding [P2], quality measurements (e.g., coordinate measurements) [P4], process monitoring [P7], industrial food prep [P3], physical therapy [P6], and running research experiments [P7]. When probed about design improvements, most participants noted that they found the current prototype comfortable, including switching between multiple grasps for fine or coarse motor control [P8]. A few participants desired further flexibility to grasp the interface, such as through additional handles [P7] or modifications that accommodate two-handed grasps for better control [P9].
Participant suggestions and interactions identified several opportunities to improve the demonstration modes. For the natural mode, additional markers (particularly on the top of the device) may be beneficial for tasks where the camera is above the tool. Additionally, P3 suggested using a camera with a wider field of view. From observing participants, we also believe that enabling participants to manually grab and adjust the camera may be beneficial to escape undesired views by the robot optimization. For teleoperation, participants suggested that haptic feedback and other means of input might make teleoperation easier, particularly given the learning curve of the SpaceMouse [P9]. While participants generally thought it was simple to switch between the different modes, they suggested future design iterations could simplify the process of attaching and detaching the tool for the natural mode (e.g., requiring only one hand or prompting the robot to stop moving).
V-B Limitations and Future Work
@@
In addition to the suggestions from our study participants, there were limitations of our setup and study that we plan to address in future work. We plan to test with several collaborative robots, different input devices for robot teleoperation, and through a wider suite of representative manufacturing tasks. Our study also only focused on perceptions of the VDI. Future iterations (with updated interface designs) will focus on characterizing the quality of data collection with the VDI modalities, investigating impacts of the multiple demonstration modalities on robot policy training, and exploring opportunities for robot to elicit targeted demonstration types based on task understanding.
V-C Conclusion
@@
We presented Versatile Demonstration Interface (VDI), an interface that attaches to existing collaborative robot platforms and enables operators to teach robots through three types of demonstrations: teleoperation, kinesthetic, and natural. Through a user study with manufacturing experts, we identified opportunities and feedback for future interfaces toward flexible Learning from Demonstration.

### v2

@@ IV-B Manufacturing Study Population
After completing the three conditions of the rolling task, the participants completed the press-fitting task. Participants were instructed about the goals of the task and shown how to complete the task without the robot. Participants were prompted that they could switch demonstration modes at any point during the task, but that they needed to use each mode to complete at least one press fitting. Following the two tasks, participants engaged in a semi-structured interview where they were asked about their preferences, design feedback, and potential use cases for our VDI implementation.
Metrics – Most critical to our evaluation objective was the final semi-structured interview. Responses from the semi-structured interview were synthesized into the key themes presented in the next section. We additionally collected task load, usability, and preference data to confirm that our combined demonstration interface is consistent with modality findings from previous studies [33, 32]. After each rolling task, participants filled out the NASA Task Load Index (TLX) [47], the Systems Usability Scale (SUS) Questionnaire [48], and the Usefulness and Ease of Use 7-point subscales of the USE Questionnaire [49]. Given our small sample size, we report only descriptive statistics for the questionnaire responses.
@@ V Results and General Discussion
Here we discuss the results of our user study and takeaways, limitations and future work, and general conclusions. All participants were generally able to complete the rolling task with each demonstration modality. In the press-fitting task, all participants were able to complete the task using the VDI tool. The questionnaire response results are shown in Figure 5 and participants ranked preferences are shown in Table II. Two participants responses are omitted from the questionnaire results due to missed survey items.
@@ V-A User Study Interactions and Implications
The natural demonstration form factor is appropriate, but robot state tracking needs improvements for industrial tasks. Consistent with previous studies, the majority of participants ranked the natural mode of VDI as their top preference and most favorable in terms of usability, usefulness, and ease of use. Regarding the human interface, we found that all participants chose to use the precision grasp. We were encouraged that several participants commented that the form factor felt familiar, similar to using a pencil [P3] or a manual tool [P1]. Despite the length of the tool (to accommodate force sensing), participants commented that the natural method was still the fastest way to accomplish the tasks [P7], would be preferred for tasks with delicate motions [P4], and that they generally felt they could perform the task better compared to the other demonstration modes [P5, P8]. Participants identified a range of manufacturing applications for natural demonstrations, including molding [P3], industrial food preparation [P3], coordinate metrology [P4], prepping and crimping wires [P5], powder coating [P7], material removal [P9], and sculpting [P9].
Participants had varied levels of success interacting with the natural mode’s camera tracking system. Figure 6 illustrates common successful participant strategies and failure cases. Many participants found success by avoiding large coarse motions and angling and rotating the tool such that the AprilTags were clearly visible at all times. Empirically, we found that this participant strategy tended to avoid local minima in the camera optimization. Less successful participants tended to exclusively focus on the task and at times ignored the camera optimization beeping indicator. Our optimization contained constraints to limit camera motion to a small envelope for participant safety. In the prototype, participants commented that the range was very limited [P6] and that it could at times be difficult to guide the robot to certain views of the task (i.e., the optimization encountered local minima). Given such limitations in range, P3 mentioned the natural mode may be most effective for 2D (e.g., tabletop) tasks where a similar camera angle can be used to track all hand tool poses.
From participant interviews and observations during the study, we believe a combination of future design modifications and participant training can further improve VDI’s natural demonstration experience. A few participants desired further flexibility to grasp the interface, such as through additional handles [P7] or modifications to accommodate two-handed grasps for better control [P9]. It would also be beneficial to add state-tracking markers on top of the human interface for scenarios where the camera is above the tool, though we assert the camera-pose optimization should try to escape such configurations as they occlude the view of the tool and propagate pose estimation error. Regarding the camera motion constraints, we believe that practical setups can consider wider ranges of motion as long as velocity is limited to minimize unsafe robot impacts with workers. Additionally, P3 suggested using a tracking camera with a wider field of view, which could aid in better state tracking with less robot motion. The main empirical limitation we encountered were situations where the camera would become stuck trying to follow the user’s motion. In these cases, participants became frustrated trying to guide the optimization out of the local minima and back to a more favorable pose. We believe that in future iterations, a critical functionality for the natural mode will be the option for the human to manually reposition the robot to a new desired view. Finally, given that several participants arrived at strategies (e.g., tilting the tool) that led to more favorable tracking, we believe that future user education and training can also help improve the overall experience.
Figure 6: Examples of successful strategies and failure cases of the natural demonstration mode. a) Several participants angled the front of the tool (with the most AprilTags) toward the robot during the task, which minimized movement of the robot camera and generally improved tool tracking. b) The robot could sometimes see the end effector, but none of the Apriltags which caused tracking failures. c) Occasionally, participants faced the top of the human interface toward the camera which did not include tracking tags. d) Some participants would stop the task and present the tool to the robot when it beeped. While this would temporarily restore state estimation, it led to later tracking issues when participants reoriented the interface to continue.
There are opportunities for each demonstration mode in manufacturing depending on user or task characteristics. Consistent with previous studies, we found that kinesthetic teaching was generally preferred over teleoperation. We also found that the usability, usefulness, and ease of use was worst for teleoperation and second best for kinesthetic teaching. Despite these general trends, participants thought that operator background might heavily influence desired demonstration modes for manufacturing floor workers. As P7 commented, “With the engineers that I work with… we have a broad range of right-out-of-school to thirty years’ experience. People would pick different methods for sure.” For example, several participants noted that teleoperation would likely be popular with young manufacturing workers, who might have more video game experience. With our experienced study population, teleoperation was not preferred.
While in some cases, operator differences may govern the choice of demonstration modality, task characteristics would be the more likely driver in manufacturing. As indicated above, the natural mode was preferred for tasks with precision requirements, localized motions (for easier robot tracking), and where it is feasible for the worker to be collocated. For kinesthetic teaching, participants imagined use in manufacturing applications where more direct control was desired, but where the robot could assist with pushing/pulling [P5], or maintaining orientation [P1,P2]. Such assistive applications would make it so ‘I didn’t have to do too much, but I had some control [P8]”. Participants thought kinesthetic demonstrations would be well-suited for tasks such as screw fittings on a small assembly line [P1], stoppering [P1], tapping holes [P2], grinding [P2], and drilling [P4]. Regarding kinesthetic interactions with VDI, we observed that participants often occluded the robot-mounted camera with their hand or arm. Thus, we believe a future implementation needs either (1) further optimization of the camera placement (e.g., further out from the robot) or (2) indications on the interface for how the interface can be grasped without occlusion.
For teleoperation, participants saw value for tasks where the demonstrator couldn’t be safely near the robot, such as environments with bio-hazards [P6] or heavy payloads [P7]. Participants also commented on the value of using teleoperation in manufacturing to scale down human input for small, precise motions [P4, P9] or to avoid the fatigue associated with more manual demonstration methods of repetitive tasks [P4]. Applications such as soldering or assembly with scaled-down motion [P3], corrosive material handling [P4], process monitoring [P7], and high-strength press fits [P9] were identified as good manufacturing applications for teleoperation.
There was limited support for manufacturing applications where switching demonstration modes would be valuable within a task. Several participants mentioned using one demonstration mode for more coarse motions (e.g., kinesthetic teaching) and one mode for more precise motions (e.g., natural or teleoperation) [P2, P3, P9]. However, most participants could not point to a manufacturing application where this behavior would be useful. Thus, we believe the primary value of VDI in manufacturing lies in the interface’s flexibility for different users or task characteristics.
V-B Study Limitations
@@
Our study only focused on perception and manufacturing use cases for the interaction modalities of the VDI. While we believe this initial study was critical to understand viability for VDI in manufacturing applications, further iteration on the physical design (following our study takeaways and design guidelines) and testing of VDI is needed. Future studies should focus on improving and characterizing the quality of VDI data collection, investigating impacts of the demonstration modalities (as well as combinations of demonstration data) on robot policy training, and exploring opportunities for robots to elicit targeted demonstration types from operators. Future studies should also test with several robots, more teleoperation input devices, through a wider suite of representative manufacturing tasks, and in domains outside of manufacturing (e.g., medical applications).
V-C Conclusion and Implications for LfD
@@
We presented Versatile Demonstration Interface (VDI), an interface that attaches to existing collaborative robot platforms and enables operators to teach robots through three types of demonstrations: teleoperation, kinesthetic, and natural. From our study with manufacturing experts, we highlighted the value of natural demonstrations with the robot-tracked instrumented tool, confirmed the utility of VDI in manufacturing settings through varied use cases, and assessed the role of VDI mode switching in manufacturing tasks. Future work will build more mature interfaces for data collection and assess VDI in robot policy learning.